stranger2:

Geometry help @jhonny9

stranger2:

jhonyy9:

so sin 60 is given and you know that there is a right angle with measure of 90 degrees how many will be the 3rd angle measure ?

jhonyy9:

together give up 180 degrees

jhonyy9:

the 3rd measure = 180 -(90+60)

jhonyy9:

3rd measure = 180-150 3rd measure = ?

jhonyy9:

are you here pls ?

jhonyy9:

@supie

snowflake0531:

i think it's good to point out first thing that this si a 30, 60, 90 triangle, since it's 1, 2, sqrt 3

stranger2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 i think it's good to point out first thing that this si a 30, 60, 90 triangle, since it's 1, 2, sqrt 3 \(\color{#0cbb34}{\text{End of Quote}}\) ohh-

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 together give up 180 degrees \(\color{#0cbb34}{\text{End of Quote}}\) can i put it on soundcloud or no?

darkknight:

going off of the fact that this is a special right triangle, u can do this pretty good with a picture. \[\sin(60)=\sqrt{3}/2\] This means the hypotenuse is 2. while teh opposite side to this angle is sqrt(3)  This is a right triangle so u can use pythagorean theorem to find that adjacent side length \[a^2+b^2=c^2\] Once you find that last side length, u can do this but some hints are to not forget that cos is adjacent/hypo and the answer choices are asking about any of the 2 angles, so its adjacent/hypo with respect to what angle is what you are looking for

darkknight:

lmk if u need more clarification